4sin inverse(x)+cos inverse(x)=pie Then find the value of x ?

Grade:12

3 Answers

Dhruvit Raithatha
6 years ago
[sin-1(x) + cos-1(x)] + 3sin-1(x) = π[π/2] + 3sin-1(x) = π3sin-1(x) = π/2sin-1(x) = π/6x = sin(π/6)x = 1/2
Kunal Jain
6 years ago
4sin-1x + cos-x =π3sin-1x +sin-1x +cos-1x =π3sin-1x +π/2 =π3sin-1x =π-π/23sin-1x = π/2sin-1x= π/6x = sin π/6x = 1/2. Ans.
umar hamid
6 years ago
4sin-1x + cos-1x = pi => 4sin-1x + (pi /2 - sin-1x) = pi [ sin-1x + cos-1x = pi/2] => 3sin-1x = pi/2 => sin-1x =pi/6 => x = sin( pi/6 ) =>x= sin 30 degree => x= 1/2 ANSWER. ........